OpenStudy (anonymous):

maria bought two plots of land for a total of $120,000.on the first plot,she made a profit of 15%. On the second she lost 10%.Her total profit was $5,500.How much did she pay for each plot.

OpenStudy (anonymous):

a + b = 120,000 --> a = 120,000 - b .15a - .10b = 5,500 .15(120,000 - b) - .10b = 5,500 18,000 - .15b - .10b = 5500 18,000 - .25b = 5500 -.25b = 5500 - 18000 -.25b = - 12,500 b = 50,000 a + b = 120,000 a + 50,000 = 120,000 a = 120,000 - 50,000 a = 70,000 check.. .15a - .10b = 5500 .15(70,000) - .10(50,000) = 5500 10,500 - 5000 = 5500 5500 = 5500 (correct) I have no idea if I did this correctly....you will need a second opinion.

OpenStudy (mathstudent55):

Plot 1: cost = x; profit 15% of x = 0.15x Plot 2: cost = y; loss = 10% of y = 0.1y Total cost: x + y = 120,000 Total profit: 0.15x - 0.1y = 5,500 Solve the two eqautions as a system of equations.

OpenStudy (mathstudent55):

@kelliegirl33 wrote "I have no idea if I did this correctly....you will need a second opinion." Don't sell yourself short. You did a great job, and your answer is correct.

OpenStudy (anonymous):

thanks....I wasn't even sure if I did it right.....thanks again for confirming

OpenStudy (mathstudent55):

Your check at the end shows you're right.

OpenStudy (anonymous):

@mathstudent55 can you also work out yours and let me see

OpenStudy (mathstudent55):

x + y = 120,000 0.15x - 0.1y = 5,500 Multiply second equation by 10 and write it under the first equation: x + y = 120,000 1.5x - y = 55,000 Add the equations together: 2.5x = 175,000 Divide both sides by 2.5: x = 70,000 Substitute x in first equation and solve for y: 70,000 + y = 120,000 Subtract 70,000 from both sides: y = 50,000 Answer: She paid $50,000 for the first plot and $70,000 for the second plot.

OpenStudy (anonymous):

THANKS SO MUCH.U GUD

OpenStudy (mathstudent55):

You're welcome